Hello, if someone could help me solve this problem/explain to me the steps, that would be very helpful. Thank you so much.
G and H are mutually exclusive events.
P(G) = 0.5
P(H) = 0.3
P(H|G) = 0.6

1. Find P(H OR G).
2. Are G and H independent or dependent events. Why?
I mostly just need help finding P (H OR G).

Answer by ikleyn(52818)

.

P(H|G) = 0.6 means that  the ratio  P(H ∩ G) / P(G) = 0.6  (by the definition of conditional probability).


Substitute here the given data, and you will get

    P(H ∩ G) / 0.5 = 0.6.


It implies  P(H ∩ G) = 0.5*0.6 = 0.3.



Finally, use the general probability formula for the union

    P(H OR G) = P(H) + P(G) - P(H ∩ G),

which is valid for any events H and G.


Substitute all given and found values into this formula, and you will get

    P(H OR G) = 0.3 + 0.5 - 0.3 = 0.5.

Solved and explained in all details.

--------------

In this site, there are two lessons for problems on conditional probability. They are
    - Conditional probability problems
    - Conditional probability problems REVISITED

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic  "Solved problems on Probability"
and  "Additional problems on Probability".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.